The width of a rectangle is 12 in. more than the length.
The area is 133 in^2.
Find the length and the width of the rectangle.

Answer by checkley77(12844)

W=L+12
LW=13
L(L+12)=133
L^2+12L-133=0
(L-7)(X+19)=0
L-7=0
L=7 IN. IS THE LENGTH.
7+12=19 IN. IS THE WIDTH.
PROOF:
7*19=133
133=133